The difference between twice Bill’s age and Carlos’s age is 26. The sum of Anna’s age, three times Bill’s
age, and Carlos’s age is 92. The total of the three ages is 52. How old is each person?

Answer by Boreal(15235)

2B-C=26
3B+A+C=92
B+C+A=52
C=2B-26
so 3B+A+2B-26=92 or 5B+A=118
and
B+2B-26+A=52 or 3B+A=78
2B=40, Bill is 20
Carlos must be 40-C=26 so Carlos is 14
Anna would be 18
18+3(20)+14 is 92 to check.